### v1

# TabSketchFM: Sketch-based Tabular Representation Learning for Data Discovery over Data Lakes

Abstract.
Enterprises have a growing need to identify relevant tables in data lakes; e.g. tables that are unionable, joinable, or subsets of each other. Tabular neural models can be helpful for such data discovery tasks. In this paper, we present TabSketchFM, a neural tabular model for data discovery over data lakes. First, we propose a novel pre-training sketch-based approach to enhance the effectiveness of data discovery techniques in neural tabular models. Second, to further finetune the pretrained model for several downstream tasks, we develop LakeBench, a collection of 8 benchmarks to help with different data discovery tasks such as finding tasks that are unionable, joinable, or subsets of each other. We then show on these finetuning tasks that TabSketchFM achieves state-of-the art performance compared to existing neural models. Third, we use these finetuned models to search for tables that are unionable, joinable, or can be subsets of each other. Our results demonstrate improvements in F1 scores for search compared to state-of-the-art techniques (even up to 70% improvement in a joinable search benchmark). Finally, we show significant transfer across datasets and tasks establishing that our model can generalize across different tasks over different data lakes.
* Joint first-authors; # Work done while at IBM Research.
PVLDB Reference Format:
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://anonymous.4open.science/r/tabsketchfm-6CDC/README.md.
1. Introduction
Enterprises store critical data in data lakes, large repositories of tabular data, for both governance and analytics (Nargesian et al., 2019). It is essential to effectively find relevant tables (e.g., joinable, unionable, subsets) within lakes for reporting, decision-making, statistical analysis, and more (Khatiwada et al., 2022; Khatiwada et al., 2023b). For instance, Unionable table search helps to augment an existing table with new rows, and to enrich analytics over it (Nargesian et al., 2018). Joinable table search is useful to run analytics that needs data in columns from multiple tables, e.g., to identify how many customers in some region purchased a product due to an email campaign. Also, identifying subsets or supersets of a table is essential, for example, when searching for potential copies of personal data to comply with the General Data Privacy Regulation of the European Union [^1] https://gdpr-info.eu. Motivated by such use cases, we focus on the problem of identifying tables from the data lakes that can be unionable, joinable, and/or subsets of each other  (Nargesian et al., 2018; Zhu et al., 2019; Khatiwada et al., 2023a). Specifically, in this work, we propose a pre-trained tabular model that could help in enhancing the effectiveness of data discovery techniques.
Central to the problem of identifying relevant tables in the data lakes is identifying similar columns in tables. However, the columns in a table could be ambiguous. For instance, a column called Age is different when it is in a table about buildings from when it is in a table about people. Furthermore, values within a column Age play a key role in determining semantics; Age of buildings in New York are unlikely to match Age of historical buildings in Egypt. Contextualizing columns could be useful in resolving such ambiguities. Language Models (LMs) are shown to be successful in contextualizing words in a sentence  (Devlin et al., 2019). Consequently, we use them to contextualize columns and use such contextualization to determine similar columns or tables.
Existing pre-trained tabular models (Deng et al., 2020; Iida et al., 2021; Wang et al., 2021a; Yin et al., 2020) based on LMs (Devlin et al., 2019) do not focus on identifying relevant tables. Instead, they focus on language-style tasks such as fact checking (Wang et al., 2021a), question answering (Herzig et al., 2020), converting natural language to SQL (Yin et al., 2020), table cell content population (Deng et al., 2020), and table metadata prediction (Iida et al., 2021; Deng et al., 2020). In addition, input to these models is the actual cell values or a row because of the focus on language style tasks. However, this has two limitations for understanding relevant tables: (a) only a small number of values can be passed in as input because of severe limits in the length of input token (e.g. 512 tokens in BERT (Devlin et al., 2019)), and this potentially limits the information available to the model to understand the input table; (b) treating numerical values in the table as text tends to lose their semantics; it maybe better to pass numerical information directly as vectors directly into the model.
To address these limitations on input length and numerical representation, we introduce a novel transformer-based tabular model—TabSketchFM—exploiting both table metadata and cell values that creates different sketches over them instead of linearizing cell values. Specifically, we abstract column cell values using data sketches, as has been used in the data discovery literature (e.g., MinHash sketches  (Zhu et al., 2016), and numerical sketches  (Santos et al., 2021)). Such sketches capture tabular features, and we input them into the neural models. Inputting numerical sketches bypasses feeding table cell values to the model, addressing the token limit issue, as well as the numerical representation of cell values.
Following a phase of pretraining with sketches, we fine tuned the models for the data discovery tasks of union, join, and finding subsets. While some in the literature (Fan et al., 2023) have used self supervision for fine tuning, self supervision applies better for some tasks such as unions compared to say joinability. To finetune models for the data discovery tasks, we developed a suite of 8 benchmarks across the 3 tasks LakeBench that we open source for use. On the fine tuning tasks, TabSketchFM outperformed other neural methods, by as much as 55% on F1. We then used the fine tuned models for search tasks, and showed that TabSketchFM outperforms the state-of-the-art neural and traditional methods by as much as 70% on F1.
In summary, our contributions are as follows:
  • A novel tabular pretrained model. We are the first to present a sketch-based tabular pretrained model, TabSketchFM, which contextualizes columns based on their data sketches and tabular metadata. The sketches help to better handle input token limits and numerical cell values. To achieve this, we modified the text-based transformer architecture to take in numerical and MinHash sketches as input. A purely sketch based approach though misses the opportunity to include table semantics from column values. Given that the inputs to the transformer architecture in our system are numeric vectors, it is difficult to combine these numeric inputs with tokens derived from cell values. We therefore combine the sketch based model embeddings with off the shelf embeddings of column values from existing pretrained sentence transformer models in a novel way. We show that this combination boosts performance on tasks.
  • Finetuning benchmarks for identifying relevant tables. We develop LakeBench, a collection of 8 benchmarks that can be used to build cross encoder models to identify relevant tables. Our benchmark consists of enterprise and open data tables, and tables generated using knowledge graphs. Across the suite, we have a mix of real benchmarks and synthetically derived ones, and a variety of fine tuning tasks such as binary classification, multi-class classification, and regression.
  • Detailed ablation study. We assess the effectiveness of different sketches for identifying different types of table relevance. Our study presents interesting insights such as Minhash sketches of column values are crucial for identifying joins, but numerical sketches are crucial for identifying subsets; once again highlighting the value of using sketches for table representation.
  • Demonstration of TabSketchFM on search. We compared the performance of TabSketchFM’s fine tuned models against neural and non-neural baselines. In the case of join search, TabSketchFM outperformed the nearest systems geared for join search by almost 70%. In addition, the model’s performance was robust against transfer across datasets and tasks.
2. Related Work
Tabular Pretrained Models. Neural models are pretrained on tabular datasets to either recover masked tokens in the table or to detect corrupt values (Deng et al., 2020; Iida et al., 2021; Wang et al., 2021a; Yin et al., 2020). Badaro et al. (2023) surveys neural models for tabular data representation and their applications. Deng et al. (2020) combined the Masked Language Model (MLM) objective from BERT (Devlin et al., 2019) with a novel Masked Entity Recovery to train TURL. Wang et al. (2021a) use MLM with novel Cell-level Cloze and Table Context Retrival for TUTA. Yin et al. (2020) used Masked Column Prediction and Cell Value Recovery for TABERT while Iida et al. (2021) repurposed ELECTRA’s objective function (Clark et al., 2020) for TABBIE. These models are finetuned and evaluated for different downstream tasks such as table metadata prediction, table content population, fact-checking, question answering, semantic parsing, and so on. For example, TABERT is evaluated on neural semantic parsing; TABBIE for column, row, and column type population tasks; and TUTA is finetuned and evaluated for cell and table type classification tasks. These downstream tasks take either a table or a table with a question as input; they do not consider the inter-table tasks such as dataset discovery, which we consider in this paper. Furthermore, they adopt serialization techniques from LMs and represent table values as sentences, whereas we represent the tables efficiently by using sketches.
Dataset Discovery Methods. Different methods such as keyword search (Cafarella et al., 2008; Brickley et al., 2019) and table search (Sarma et al., 2012; Lehmberg and Bizer, 2017; Fernandez et al., 2018) have been used for the dataset discovery over data lakes. Nargesian et al. (2018) (TUS) presented finding top-k data lake tables that are unionable with a query table. They determine the column unionability using three statistical tests based on value overlap, semantic overlap, and word embedding similarity, and aggregate them to infer table unionability. In $D^{3}L$,  Bogatu et al. (2020) adopted value overlap and word embedding measures from TUS and added three additional measures to measure column unionability: numerical column distributions, column header similarity, and regular expression matching. Khatiwada et al. (2023a) introduced SANTOS which considers the relationships between the column pairs in both query and data lake tables. Fan et al. (2023) developed a contrastive-learning-based table union search technique called Starmie that captures the context of the whole table in the form of column embeddings. Recently, Hu et al. (2023) presented AutoTUS which searches for unionable tables by contextualizing entire tables with embeddings of the relationship between the column pairs. Furthermore, there are other table search techniques that focus on finding joinable tables, i.e., given a query table marked with a query column, find the data lake tables that contain a column with overlapping values with the query column. For instance, Zhu et al. (2016) defines joinable table search as a problem of finding data lake tables that have a column having high set containment with the query column. They define a novel LSH Ensemble index that approximates set containment computation. Zhu et al. (2019) developed JOSIE that searches for top-k joinable tables based on exact set containment. Dong et al. (2021) presented PEXESO, which efficiently searches for the top-k semantically joinable tables. Along with a join on exact values, they also consider fuzzy join, i.e., join on abbreviations and synonyms. Dong et al. (2023) introduced DeepJoin, a deep-learning-based technique that also searches for semantically joinable tables. Recently, Cong et al. (2023) also proposed WarpGate, a system for searching for top-k semantically joinable tables by embedding each column to vector space using pre-trained embeddings. They index column embeddings using Locally Sensitive Hashing (Gionis et al., 1999) for efficient search. In line with these data discovery methods, we also address the task of discovering tables from the data lakes that can be unionable, joinable, and/or subsets of each other. Notably, we propose a single pretrained neural model that can be specialized for any of the above tasks. The advantage of our approach over task specific deep learning based approaches (Fan et al., 2023; Dong et al., 2023) is that, as in natural language processing, much smaller datasets are sufficient to finetune new data discovery tasks leveraging the existing tabular knowledge in the pretrained model.
Benchmarks. Several tabular benchmarks exist in the literature  (Badaro et al., 2023; Hulsebos et al., 2023). Koutras et al. (2021) created benchmarks for schema matching (Rahm and Bernstein, 2001) using open data tables and evaluated the existing matching techniques over them. Mudgal et al. (2018) open-sourced 13 datasets such as the DBLP-Google Scholar author dataset and Walmart-Amazon product datasets for entity matching task (Konda et al., 2016). The WDC Web Table corpus (Lehmberg et al., 2016), containing around 233 million web tables extracted from different web pages, has been widely used for diverse tasks such as Question Answering (Herzig et al., 2020; Liu et al., 2022), Semantic Parsing (Yu et al., 2021; Liu et al., 2022; Yin et al., 2020), Table Retrieval (Wang et al., 2021b), Table Metadata Prediction (Wang et al., 2021a; Deng et al., 2020; Suhara et al., 2022) and more (Zhang and Balog, 2020). Recently, Efthymiou et al. (2022) used knowledge graphs to create tabular benchmarks for column type prediction and column-to-column binary relationship prediction tasks. The VizNet Benchmark (Hu et al., 2019) has been used to evaluate column type prediction tasks (Zhang et al., 2020; Hulsebos et al., 2019). But the most closely related benchmarks to our work are TUS (Nargesian et al., 2018), SANTOS (Khatiwada et al., 2023a), and UGEN-V1 (Pal et al., 2023) released for table union search. These search benchmarks are, however, not designed for training neural models. To create such benchmarks, we adapt the existing public datasets for pairwise classification and regression tasks. GitTables (Hulsebos et al., 2023) is a recent large table corpus in the literature. However, they contain web tables that are small entity tables with fewer rows and columns than enterprise tables and do not represent enterprise data lake scenarios (Zhu et al., 2019). We use open and enterprise datasets to create our benchmarks where the tables have 10s of columns and 1000s of rows and their cell values are often numbers or cryptic codes.
3. TabSketchFM
Figure 1. Incorporation of sketches into transformer architecture. MinHashes are computed over the whole cell value, and tokens of a cell value for a given column. MinHashes are also computed over the entire row, treated as a string. For a given column token (e.g., Price in this figure), hidden states for each encoding are summed after passing through an embedding or a linear layer before it gets sent to the BERT encoder layers. Table description tokens are summed with MinHashes over the entire row.
In this section, we detail the architecture of TabSketchFM. Like existing tabular models (Wang et al., 2021a; Herzig et al., 2020), we pretrain a transformer-based BERT model (Devlin et al., 2019), but change the architecture to accommodate numerical inputs. The goal of building the pretrained model is to create cross encoders that can be used for ranking in data discovery using significantly smaller sized labeled datasets.
The first significant departure from other tabular representation models is in how we represent the table as input. While other models linearize table cells as text, we create a sketch for each column which are numerical vectors that need to be translated into a form that can be consumed by the transformer architecture. The sketches capture different features of the columns. Figure 1 shows an overview of how the different sketches we describe below are fed to the self-attention layers. Specifically, we create the following sketches for each column.
1. Numerical sketches. For each column, we record the number of NaNs and the number of unique values (normalized by the number of rows in the table). For string columns, we also record cell width in bytes, since that can be an important determinant in governing if the column is likely to be a join column. For instance, very long strings are unlikely to be join candidates. When possible, we convert date columns into timestamps and then treat them as numeric columns. For numeric columns, we additionally compute a percentile sketch, and add the mean, the standard deviation, and minimum and maximum values. All these features are encoded as individual elements of a single input vector we call the numerical sketch.
2. MinHash sketches. For each column, we compute a MinHash signature of cell values. [^2] In our implementation, we use datasketch library (Zhu and Markovtsev, 2017) to compute MinHash. As shown in Fig. 1, the cell value of Austria Vienna is used as a single element in the set passed to the MinHash computation. [^3] While it does not make sense to convert float or integer columns into MinHash signatures, we do so because it is often difficult to tell if a column is truly a float or an integer, or it really is a categorical value. For string columns, we also compute another MinHash signature for the tokens within the column. The rationale for this second MinHash is that the tokens can sometimes capture its semantics (e.g., if a token street appears in two different columns, they maybe both about address, which is useful if one considers a union of two tables that do not necessarily share cell values). As shown in Fig. 1, the cell values of Austria and Salzburg get tokenized and used as two different elements of the set passed to the MinHash computation. We concatenate both MinHash sketches into a single input vector for string columns. For numerical and date columns, only the MinHash for the cell values is included in the input vector.
3. Content snapshot. Row-based information could be crucial in understanding the table semantics for data discovery (Khatiwada et al., 2023a). We concatenate values within each row to form a string and create MinHash signature for each of them. For instance, we concatenate the last row of table in Fig. 1, surrounded by a green rectangle, into "Quarterly Austria Salzburg 800,000 28/03/23" and hash it as another minhash style sketch.
Now that we have these sketches, which are arrays of numbers, a key question is how to adapt BERT’s architecture to consume these numeric values. Transformer models take sentences as input, and internally break it into different inputs such as the tokens of a sentence [^4] The BERT tokenizer sometimes breaks a single word into multiple tokens such that unknown words in its vocabulary are handled correctly., the position of tokens in the sentence, and a scalar type indicating if the token belongs to the first sentence or the second. We leverage this mechanism to pass the table-specific scalar inputs as shown in  Fig. 1 to the BERT Embedding Layers. For vector inputs (e.g., numerical sketch, MinHash sketch), we expand the notion of BERT Embeddings to add some linear layers, as described below.
TabSketchFM’s specialized layers for processing tabular input are shown in the light gray box, middle-right of Fig. 1. The token Price for the Price column is used as an example for illustrative purposes. All other tokens belonging to the table description and other columns would undergo a similar process to create a single embedding per token.
1. Token Embedding Layer. The embedding layer of the BERT uncased model (Devlin et al., 2019) contains numerical embeddings for each word in its vocabulary based on extensive training with a large corpus, such that, for instance, the embedding of Price might be closely associated with that of House because they co-occur often in text. Our model’s token embedding layer is initialized with the weights of the BERT model’s embedding layer to leverage such information present in natural language, but the weights are allowed to be changed further while pretraining to include co-occurrence information in table-specific corpora. As shown in Fig. 1, the token Price would be mapped from its position in the vocabulary of the model (also pre-populated with BERT’s vocabulary) into its hidden state using weights from this token embedding layer.
2. Token Position Embedding Layer. Each column is analogous to a sentence in a paragraph. Hence, we re-purpose token positional embeddings to reflect a token’s position within a column name. For instance, as shown by the light gray scalar input in Fig. 1, the token position of a token Area is 1 because it is the second token in the cell (after Reference).
3. Column Position Embedding Layer. Columns themselves have positions and are encoded through a new embedding layer which can range from 1 to the total number of columns, [^5] Position 0 is reserved for any tokens in the table description. as shown by the white scalars in Fig. 1. The rationale for including column positions is that they sometimes do have semantics; e.g., a street address next to the city, followed by a zipcode. Of course table semantics do not change as a function of column order; nevertheless, we included position in case it helps the model understand column semantics, with the assumption that the attentional mechanism would disregard order if necessary.
4. Column Type Embedding Layer. Column types of string, date, integer, or float are encoded through another embedding. In order to determine column type, we made a best-case effort to parse the first 10 values of each column as dates, integers, or floats and defaulted to string if we could not convert them. In mixed-type columns, this can yield poor results, but at the very worst, at least one of the types was assigned to these columns. Column types are shown in Fig. 1 in orange.
5. MinHash Sketch Linear Layer. MinHash sketches and the content snapshots were passed through a linear layer to encode them into the same hidden state dimensions of the existing transformer layers in Bert.
6. Numerical Sketch Linear Layer. Numerical sketches are similarly passed through a linear layer to encode them.
As shown in Fig. 1, for each column token, we combined the hidden states of its token embedding, token position embedding, column position embedding, column type embedding, MinHash sketch embedding, and numerical sketch embedding using summation. This summation is similar to the summation of token embeddings and position embeddings in the original BERT model. Content snapshot signatures are summed with table description tokens because the content snapshot describes the entire table’s content. For instance, for the tokens Residential, Property and Prices, the content snapshot is passed into the linear layers to create a hidden state that is combined with the other embeddings for those tokens. Once each token has been summed with its other sketches, positional and type encoding, it is passed to the rest of the encoder and attention layers from BERT. The attention layers are crucial for contextualizing the embeddings based on what other columns, values, and table descriptions exist in the table. As shown in the figure for Price, the attention mechanism learns $N^{2}$ attention weights between all $N$ tokens for a given table. This step will consider Price in relation to all other columns as well as table descriptions. We used the usual 12-layer BERT encoder model with self attention from HuggingFace [^6] https://huggingface.co/docs/transformers/model_doc/bert once we have unified the hidden states from the embedding layers and the linear layers.
4. Pretraining
Dataset. Existing works generally use non-public data (Badaro et al., 2023) or web tables (Lehmberg et al., 2016) to pretrain the models. Unlike enterprise data, web tables often have few rows and columns to ease human consumption (Cafarella et al., 2008) and they focus on entities popular on the web (e.g., countries). Their performance may not generalize to enterprise data, where the tables have a large number of rows, the entities are often domain-specific, contain cryptic code words, and have a lot of numerical information. Therefore, we created a de-duplicated pretraining dataset of $197,254$ enterprise-like tables (CSVs) from CKAN and Socrata that are legally cleared to have the open licenses. This pretrain dataset contains $2234.5$ rows and $35.8$ columns in each table on average. About $66\%$ of columns were non-string, resembling an enterprise datalake.
Data Augmentation. In the image processing literature (Acharya et al., 2006), different transformations are applied over an original image to generate new image samples for training. This makes the trained model robust on different variants of the same image. We want the model to be robust to the order of columns such that shuffling columns in a table does not impact its semantics. So, we created three different versions of the table, by changing the column order, which in turn, changed the table’s content snapshot (see Section 3). This increases the total number of pretraining tables to 290,948 and for each of them, we create signatures as explained in Section 3.
Method. We employ the standard Mask Language Modeling (MLM) as the pretraining objective for our model, as in most text-based language models. First, a few tokens are randomly sampled according to the masked language probability. Then they are substituted with a [MASK] token to create an input sequence. The input sequence is fed to the model, and the model is asked to predict the substituted words. This can be viewed as a classification task where the original word is the label and the model’s entire vocabulary is the set of possible labels. Cross-entropy loss is computed for each predicted word in the input text that was masked.
(1) $\frac{-1}{N}(\sum_{n=1}^{N}y_{i}*log(\hat{y_{i}}))$
Equation 1 defines cross-entropy loss, where $y_{i}$ is the class labels in the ground truth, $\hat{y_{i}}$ is the predicted class label and N is the number of training examples in the train set.
Figure 2. A combination of whole column masking, and MLM probability masking to generate up to 5 examples. The box on the left hand side is the table description and the one on the right hand side is the column names
In our scenario, for each table, we mask a single column name such that all tokens corresponding to that column are masked. This is analogous to the natural language literature’s whole word masking where all tokens corresponding to a word are masked. In addition, we used MLM probability to mask the tokens in the table description as well. As shown in Fig. 2, we generate several samples from a single table. Note that each masking gives us an example. For small tables with less than $5$ columns, we mask each of its columns one after another. However, if a table has a large number of columns, this produces a lot of examples for the same table, over-representing it. So, for large tables with more than $5$ columns, we randomly select five columns and mask them. Following this strategy over the augmented pretraining dataset, we get $730,553$ examples in training, $54,430$ examples in validation, and $58,141$ examples in test sets. We provide the experimental setup for pretraining in Section 6.
5. Finetuning Datasets
Table 1. Cardinality of all the datasets in LakeBench.
Benchmark Task # Tables Avg. Rows Avg. Cols # Table Pairs Data type distribution (%)
Train Test Valid String Int. Float Date
TUS-SANTOS Binary Classification 1127 4285.17 13.04 16592 3566 3552 77.94 8.62 7.51 5.93
Wiki Union Binary Classification 40752 51.05 2.66 301108 37638 37638 57.97 14.38 24.25 3.4
ECB Union Regression 4226 292.47 36.3 15344 1910 1906 47.72 14.05 36.31 1.92
Wiki Jaccard Regression 8489 47.3 2.8 12703 1412 1572 57.5 15.66 19.76 7.07
Wiki Containment Regression 10318 47.15 2.79 21007 2343 2593 57.26 15.26 20.58 6.9
Spider-OpenData Binary Classification 10730 1208.87 9.09 5146 742 1474 42.22 18.51 32.62 6.66
ECB Join Multi-label Clasification 74 8772.24 34.47 1780 222 223 52.14 7.8 37.79 2.27
CKAN Subset Binary Classification 36545 1832.58 25.37 24562 2993 3010 31.75 17.53 46.14 4.58
After pretraining, we need finetuning to adapt the pretrained model to particular tasks of identifying relevant tables such as unionability, joinability, or subset. Existing finetuning datasets target other tasks such as table metadata prediction (Deng et al., 2020), question answering over tables (Yin et al., 2020), and so on, and none focus on relevant table understanding. Therefore, we create LakeBench, a collection of finetuning datasets to identify relavant table pairs using multiple data sources such as open government data from CKAN and Socrata, economic data from the European Central Bank (Eurosystem Data, 2023), Spider (Yu et al., 2018), and synthesized data from large knowledge graphs such as Wikidata (Vrandečić and Krötzsch, 2014). Crucially, we ensure that if a table is a part of the pretraining dataset, we never use it as a part of any finetuning dataset.
LakeBench covers Unionability (Section 5.1), Joinability (Section 5.2) and Subset (Section 5.3) tasks, and a wide range of problem types: binary classification, regression, and multi-label classification. Each benchmark comes with train, test, and validation sets where each example contains a table pair and its label. Table 1 summarizes the tasks and dataset cardinalities.
5.1. Unionability
Two tables A and B are unionable if a subset of their columns are unionable (Nargesian et al., 2018), and fully unionable if all their columns are unionable. A column $c_{1}$ from Table A is considered unionable with a column $c_{2}$ from Table B if they contain values that are drawn from the same domain. While overlap between values in the columns can indicate that they belong to the same domain, columns with few or no overlapping values may also map to the same semantic type or concept. For instance, two columns titled “movie" and “film" might have different movie names in them with no overlap, but they can still be unionable because they map to the same semantic type. Generally, we take a union of a new table with an existing table to expand it vertically (i.e., add more rows) (Khatiwada et al., 2023a). LakeBench contains three datasets for the unionability task: one adapted from existing table search benchmarks (Khatiwada et al., 2023a; Nargesian et al., 2018), one synthesized from Wikidata, and one from ECB data.
1. TUS-SANTOS
Nargesian et al. (2018) created the TUS Benchmark to evaluate the performance of table union search methods over an open data lake. They use $32$ large seed tables from Canada’s Open Government Data Portal (of Canada, 2023) and U.S. Government’s Open Data (Government, 2023). The seed tables are selected from unique domains such that they are not unionable with each other. Then, they generate $5000$ smaller tables by randomly sampling the rows and columns from each seed. The smaller tables generated from the same seed are unionable to each other and the tables from different seeds are not unionable. Khatiwada et al. (2023a) created SANTOS Benchmark by adapting the same benchmark creation technique (Nargesian et al., 2018), but preserving the context and the key entity column during the split. For example, if a seed table contains columns (“person", “birthplace", and “birth year"), the TUS approach might result in a table with (“birthplace" and “birth year"), while the SANTOS approach ensures that the “person" column is preserved (Khatiwada et al., 2023a). These benchmarks were constructed for a search task where the requirement is to retrieve unionable tables from a data lake given a query table. To curate a benchmark for finetuning the pretrained model for the table unionability task, we sample unionable and non-unionable table pairs from TUS and SANTOS and create a binary classification dataset.
2. Wiki Union
To curate a larger dataset along with ground truth unionability assessment labels, we use the process depicted in Fig. 3. We take a structured knowledge base (KB) with a SPARQL (Sparql data, 2023) endpoint, and create a collection of tables describing entities in the input knowledge base. The process starts with creating a “profile" of the ontology of the KB. The profile consists of a list of all the classes, the properties of each class, and statistics about the number of instances of each class, the number of values for each property, as well as the data types. This profile is then used to specify a configuration that determines the characteristics of the resulting tabular data lake. The characteristics include: domain (i.e., types of entities), inclusion and prevalence of different data types (e.g., numerical or categorical), the minimum and maximum number of rows and columns in tables, the prevalence of tables that have the same schema, the prevalence of ambiguous entity labels, the amount of noise introduced (if any), the prevalence of tables with null values, and the maximum number of tables about the same entity type. The specification is then used to generate a raw collection of tables. The result is a possibly very large collection of tables, along with ground truth mappings to the KB.
Figure 3. Using a knowledge base to curate a tabular data collection for table union and join benchmarks
Figure 4. Sample subsets of four tables from our Wikidata-based table union and join benchmark (top), along with ground truth labels and mappings (bottom)
We use Wikidata (Vrandečić and Krötzsch, 2014) as the source KB and curate a tabular data collection. The resulting collection consists of large tables with mostly numerical columns, to resemble tabular data found in enterprise data lakes. Figure 4 shows example tables from this collection. The first column in each table contains the labels of entities in the KB. Every other column represents either a relation or a property of the entity. Relation columns contain labels of entities while property columns contain literal values. Some tables may contain null values in some of their rows and one of their columns. The configuration consisted of a subset of Wikidata classes along with only their numerical properties, in addition to a random subset of classes along with all their relations and properties. The configuration specified between 2 and 8 columns, and between 20 and 120 rows. As the classes in Wikidata are quite large and may result in a large number of tables and an unbalanced collection, we limit the number of rows in all the tables for each class to 3,000, and limit the number of tables with the same schema to 20. The raw collection consists of $46,521$ tables, with $3,157,781$ mappings of cell values to $1,317,724$ unique entities, $72,458$ property mappings, and $53,087$ column to concept mappings.
To curate the Wiki Union dataset, we designate two tables as fully unionable if they are about the same concept and all their columns map to the same properties in the KB as per the ground truth mappings. The first two tables in Fig. 4 are such fully unionable tables. We create two kinds of negative labels: a) tables in which columns map to the same properties but are about different entities; and b) tables with the same number of columns but not all of their columns map to the same properties. Tables QCXMIM62QXN0.csv and GSW089XWBCBS.csv in Fig. 4 are negative examples, with both tables having columns that map to area (P2046) and population (P1082), but the tables are about different types of entities (as shown in table cta_gt.csv) and not unionable.
3. ECB Union
The European Central Bank (ECB) distributes detailed information about the economy across the European Union (Eurosystem Data, 2023), organized into $74$ datasets, each dedicated to a different topic. Each dataset contains a set of time series tables, with varying features from a given dimension. Figure 5(a) shows an example from two datasets that are related: residential property prices (RESR) with over 500 tables, and residential property valuation (RPV) with 232 tables. Each table specifies the values for each dimension in the dataset, for instance, AT stands for Austria, AT130 stands for Vienna, EC stands for everything except Vienna, etc. For the ECB union benchmark, we created a regression task by pairing tables such as the table for Vienna, new flats as shown in the figure with the table for outside Vienna, new flats and the table for outside Vienna, all flats, with the assumption that the plausibility of the union depends on the number of changed values across the dimensions. Therefore, we construct ECB Union benchmark by ranking each pair by how many dimensions differ, which varies from 1 to 12 in our benchmark dataset. This desired ranking forms a regression task.
Figure 5. (a) Example Datasets from ECB (b) Spider-OpenData Join benchmark data generation: Once a valid join column is selected, the table is sorted based on the join column, and then divided into four quadrants to identify positive (horizontal) and negative (diagonal) examples.
5.2. Joinability
We consider a column from Table A as joinable with a column from Table B if the two columns map to the same semantic type and have overlapping values. We join a new table to an existing table to expand it horizontally (i.e. to add more features or columns) (Zhu et al., 2019). LakeBench contains four datasets for the joinability task; two created from Wikidata, another which was designed from Spider and CKAN/Socrata; and the last reflects existing joins from ECB.
1. & 2. Wiki Jaccard and Wiki Containment
We use the same corpus of Wikidata tables we used for Wiki Union to derive a benchmark for table joins. For this benchmark, we use the cell-entity (CE) mappings in the ground truth mappings and assign joinability scores to pairs of columns in the collection. The score is either the Jaccard similarity (size of intersection over the size of the union) across sets of CE mappings, or the minimum containment ratio across the sets of CE mappings which indicates an overlap in the entities in those columns and so a potential for joining. Tables T1GB2LNKTOXP.csv and SW4T7EFEI2DK.csv in Fig. 4 are examples of such tables in our benchmark, because they share a number of values in their first column that map to the same knowledge base entity. These two tasks are modeled as regression tasks.
3. Spider-OpenData Join
To create this benchmark, we used two data sources, Spider (Yu et al., 2018) and CKAN/Socrata open government data. Spider is a large-scale human-annotated text-to-SQL dataset annotated by 11 Yale students. It comes with 10K questions and 200 databases with multiple tables covering 138 different domains. Within each database, joinability is clearly identified via primary/foreign key relationships. Due to the relatively smaller number of tables per database, we were able to generate only a small number of join examples.
To ensure that we have enough samples for training and testing various models, we also used CKAN (CKAN data, 2023) and Socrata (Socrata data, 2023) open government data. However, this dataset did not come with manually annotated joins, so in the spirit of TUS (Nargesian et al., 2018) and SANTOS (Khatiwada et al., 2023a), we synthesized joins. Figure 5(b) illustrates this benchmark creation. For every table with at least 2 columns, 1) we select a join column at random from the set of columns with mostly unique values that are not floats, 2) we sort the table based on the join column, and 3) we divide the table around the join column into four quadrants. The top two quadrants and the bottom two quadrants are good candidates for positive joinable tables. Adjacent quadrants share the same join column and hence it is considered a true positive join. To create negative examples, we pair quadrants on the diagonal to form two pairs of examples, (top left, bottom right) and (bottom left, top right). We also ensure that these negative examples are true negatives and they do not share any values across the join column.
Figure 6. Pair creation for $S_{1}$ from CKAN/Socrata. To ensure positive and negative subset pairs differ by the same amount in the number of rows, $S_{1}$ is paired with the union of $S_{1}$, $S_{2}$, $S_{3}$ for a positive pair and $S_{2}$, $S_{3}$, $S_{4}$ for a negative pair.
4. ECB Join
As shown in Figure 5(a), 74 datasets in ECB have several shared dimensions (a total of 56 dimensions in all, with some tables sharing as many as 18 dimensions). Many of those datasets such as RESR (residential property prices) and RPV (residential property valuations) are related on multiple dimensions. For the join task, we collapsed across all tables within a given dataset into a single table, just so we could evaluate joins efficiently. Many tables are therefore large; the largest one has over 20 million rows. For each pair of tables, we computed joins on all shared dimensions to see if the result returned any rows. If it did, we recorded the dimensions on which the join was possible to model it as a multi-label classification problem. If the tables shared dimensions but a join resulted in no rows, this was recorded as another label (i.e., no joins are possible). The total dataset had 1780 pairs for training, 222 for valid and 223 for test. This multi-attribute join benchmark is modeled as a multiclass classification task. Although this is a relatively small dataset for finetuning, we include it as it specifies multi-attribute joins on very large, realistic tables, which is difficult to construct synthetically. It is also a measure of how much data is actually needed to finetune a model.
Figure 7. Example Table Subsets from CKAN Subset
5.3. Table Subsets
Table A is a subset of Table B if the rows of Table A are contained in B. The subset task is a useful one for governance as it is often useful to trace the provenance of data across tables in a data lake, and to find potential copies of a table (with some modifications). Table subset is defined as a binary classification task. We provided the models with the column names in the table, which meant that positive and negative examples had the exact same schema, but differed in values. In LakeBench, we provide a subset benchmark created using tables from CKAN/Socrata.
CKAN Subset
The subset problem becomes challenging when the schemas of the positive and negative examples are exactly the same. But random pairs of tables in the CKAN/Socrata data are most likely to have different schemas. So, we adopted a strategy shown in  Fig. 6. Each table greater than 100 rows was partitioned into 4 equal subsets, $S_{1-4}$. Each subset $S_{i}$ was paired with a table composed of $S_{i}$ and with two other subsets drawn randomly (e.g., $S_{2}$,$S_{3}$ in Fig. 6) for a positive example (shown in green), and paired with a table composed of all other subsets (i.e., $S_{k\neq i}$, shown in red) for a negative example. With this strategy, the number of rows differential for positive and negative pairs could not be used as a signal for whether a pair of tables was a subset or not. Much of the CKAN/Socrata data is de-normalized (Usta and Salihoglu, 2023), so it contains many repeating values in certain columns (e.g. name of a department). This means that negative examples often overlap in some of the columns, which makes the subset problem more difficult. Figure 7 shows example table subsets from CKAN Subset.
6. Experiments
Now we empirically address the following research questions:
  1. Q1.
    How does sketch-based TabSketchFM compare to existing cell-based models for building cross encoders for the tasks in LakeBench?
  2. Q2.
    How do different sketches of TabSketchFM impact different tasks?
  3. Q3.
    How can one use the cross encoders built from TabSketchFM for search over data lakes?
Table 2. Performance (avg $\pm$ stdev) of TabSketchFM vs. other baseline models on LakeBench averaged across five random seeds. Best performance is highlighted in bold, and the second best is underlined.
Tasks Vanilla BERT TAPAS TABBIE TUTA TaBERT TabSketchFM (ours)
TUS-SANTOS (F1) $\boldsymbol{0.99\pm 0.00}$ $0.34\pm 0.00$ $\underline{0.75\pm 0.22}$ $\boldsymbol{0.99\pm 0.00}$ $\boldsymbol{0.99\pm 0.00}$ $\boldsymbol{0.99\pm 0.00}$
Wiki Union (F1) $0.33\pm 0.00$ $0.41\pm 0.00$ $0.64\pm 0.12$ $0.33\pm 0.00$ $\boldsymbol{0.97\pm 0.00}$ $\underline{0.94\pm 0.00}$
ECB Union (R2) $0.03\pm 0.03$ $-0.01\pm 0.01$ $0.02\pm 0.01$ $\underline{0.87\pm 0.01}$ $0.35\pm 0.06$ $\boldsymbol{0.90\pm 0.03}$
Wiki Jaccard (R2) $00.00\pm 0$ 0.00 $-0.03\pm 0.03$ $0.25\pm 0.02$ $\underline{0.43\pm 0.02}$ $0.33\pm 0.03$ $\boldsymbol{0.58\pm 0.03}$
Wiki Containment (R2) $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $0.21\pm 0.01$ $\underline{0.35\pm 0.03}$ $0.30\pm 0.03$ $\boldsymbol{0.58\pm 0.02}$
Spider-OpenData (F1) $0.71\pm 0.06$ $0.65\pm 0.00$ $0.57\pm 0.00$ $0.76\pm 0.12$ $\boldsymbol{0.87\pm 0.02}$ $\underline{0.83\pm 0.01}$
ECB Join (F1) $0.63\pm 0.04$ $0.40\pm 0.00$ $0.42\pm 0.03$ $\underline{0.81\pm 0.01}$ $0.79\pm 0.03$ $\boldsymbol{0.86\pm 0.01}$
CKAN Subset (F1) $0.43\pm 0.00$ $0.43\pm 0.00$ $0.43\pm 0.00$ $0.43\pm 0.00$ $\underline{0.43\pm 0.00}$ $\boldsymbol{0.98\pm 0.00}$
We pretrain TabSketchFM using $4$ A100 $40$GB GPUs for $2$ days, until the model converged. Our pretrained model contains $118M$ parameters, similar to other row-based models in our experiments. We use patience of $5$, i.e., we consider the model as having converged if the validation loss does not decrease for more than $5$ epochs. As studied in the literature (Zhou et al., 2020), this not only reduces training time but also helps us avoid overfitting. For finetuning the cross encoders, for all tasks, we use $1$ A100 $40$GB GPU. Most finetuning tasks converged within $6$ hours using the same 5 epoch patience as pretraining except for the Wiki Union benchmark which took 24 hours. We use cross-entropy loss for classification tasks, mean squared error for regression tasks, and binary cross-entropy with logit loss for multi-class classification tasks. For each finetuning task, we create a cross-encoder (Du et al., 2022), which evaluates if a pair of tables can be joined, unioned or are subsets of each other.
6.1. TabSketchFM on LakeBench Tasks
We evaluate our finetuned cross-encoders and the existing value-based tabular foundational models for the LakeBench unionability, joinability, and subset tasks.
6.1.1. Baselines
We compare TabSketchFM against four publicly available tabular foundational models: TUTA, TaBERT, TABBIE, and TAPAS. We adapted these models for Lakebench data discovery tasks by building a dual encoder architecture. Each encoder represents the pretrained model with shared parameters; and the same model is used to encode each element of a table pair. The embeddings from the last layer of the encoders were concatenated and passed through a two-layered MLP (multi-layered perceptron). We used this architecture as opposed to the cross-encoder architecture we used for TabSketchFM because in general it is unclear how to feed two different tables with different row sizes into a cross-encoder. It gets even more complicated when one considers models such as TUTA that model hierarchical information in the table as a tree. While we were able to finetune the TUTA and TaBERT code in this dual encoder architecture, the code of TAPAS and TABBIE were less amenable to finetuning. Hence, for TAPAS and TABBIE, we froze their pretrained models while finetuning, but allowed the two layers above the model to learn the weights. In this case, the pretrained model can be seen as producing an embedding that is then fed into two layered network that tunes for a specific task. We will make the code for all the baselines available. Below, we provide details on model-specific implementation issues.
  • TaBERT provides two types of embeddings for a table: context embeddings corresponding to the tokens in the context, and column embeddings corresponding to each column in the table (Yin et al., 2020). For each table, we compute both embeddings for top $10,000$ rows. In Yin et al. (2020), a content snapshot approach is described which creates synthetic rows based on the n-gram matches of the input question with the cell values. We could not experiment with this feature, since 1. this technique is not part of the authors’ provided source code [^7] https://github.com/facebookresearch/TaBERT and 2. LakeBench tasks do not contain a question. So, we compute these embeddings for each table, mean-pooled over the context tokens and tabular columns respectively. The final embedding for the table pair is computed as the concatenation of context and column embeddings obtained for each table.
  • TUTA creates a tree-based attention matrix over all the tokens in the table, which has significant memory requirements. To overcome this, we obtain the top $256$ rows and columns of the table, and the first $256$ tokens of the table sequence.
  • TAPAS requires a natural language query and a table as input (Herzig et al., 2020). For our dataset discovery tasks, we sent an empty string as a natural language query and use $512$ serialized row based version of the table as the embedding.
  • TABBIE provides row and column embeddings of the table (Iida et al., 2021). For our work, we obtained the row embeddings for each row in the table. Following the original work, we use the first 30 rows and 20 columns of the table. These row embeddings are combined using the mean operation and the resulting vector is frozen as table embedding.
We also added a naive baseline, Vanilla BERT, to examine to what extent LakeBench tasks can be performed based on column headers alone; this provides a measure of the difficulty of a given task. For Vanilla BERT, column headers were provided for the two tables as two sentences, and the BERT model was finetuned on the specific LakeBench task as were other models.
6.1.2. Results
Table 2 compares the performance of TabSketchFM with baseline models on LakeBench. For regression tasks, we report R2 statistics, and for (binary and multiclass) classification tasks, we report a weighted F1 score to handle skew in classes [^8] Our metrics were implemented from scikit-learn: https://scikit-learn.org/stable/modules/model_evaluation.html. The best performance is highlighted in bold and the second best is underlined. All the results presented are aggregated over 5 random seeds. We see that on five of the eight tasks, TabSketchFM outperformed all baselines, performed as well as the best methods on one, and performed second best on the remaining two.
The binary classification adaptation of the TUS-SANTOS benchmark, was too easy and could be solved on the basis of column headers alone [^9] We added column headers for this benchmark, unlike in the literature, but removing headers made little difference to the results of tabular models., as evident from the performance of the Vanilla BERT model. Nevertheless, TAPAS and TABBIE perform relatively poorly on it, pointing to the importance of including the Vanilla BERT as a baseline. On the CKAN Subset benchmark, we find that the performance of most models was comparable to random guessing since the column headers were exactly the same, and most systems (except TaBERT) did not have a view of the entire dataset. On Wiki Union and on Spider-OpenData, TaBERT outperformed TabSketchFM. Even so, TabSketchFM achieved the second best performance on these two tasks.
We also performed an error analysis on table pairs that TabSketchFM got wrong on Wiki-Union, Spider open data and CKAN. In all these cases, mean hamming distances of MinHash sketches for positive pairs were equivalent to those from negative pairs; suggesting that the sketches alone did not contain sufficient information to discriminate those examples. TaBERT’s superior performance on the Wiki Union case, combined with the fact that lack of discrimination of MinHashes provides a clue to at least one weakness in the current version of TabSketchFM, which is that we do not consider any values from the column at all. As shown in Figure 4, two columns may be unionable when their values map to the same domain (e.g., different municipalities in Slovakia), with minimal overlap in actual values. In such cases, a system such as TaBERT can still determine that the cell values are similar, and belong to the same semantic type, whereas TabSketchFM cannot. Incorporating cell values is something we plan to address in future work.
In summary, the answer to  Q1 is that sketches-based TabSketchFM is in general superior to value-based methods for supervised data discovery in LakeBench. Note that this evaluation does not directly speak to performance on search tasks in data discovery; but it does evaluate the usefulness of sketches for building a good cross encoder. We turn next to investigating how different sketch types help specific tasks.
6.2. Effectiveness of sketch types on task types
To explore the relevance of the sketches and content snapshots for the different data discovery tasks, we conduct an ablation study where we use only one type of sketch. The results on LakeBench tasks with the random seed 0 are shown in Table 3. In  Table 4 we present another ablation study where we remove one sketch and retained the rest. Not surprisingly, MinHash sketches are crucial for join tasks, where MinHash alone seems to dominate overall performance of TabSketchFM. Subset selection clearly relies on the overlap of the data distribution; numerical sketches alone match the performance of TabSketchFM. Given that almost 69% of the columns were numerical in the subset benchmark, this outcome might be expected. The effect of content snapshot was less clear. Looking at Table 4, it is evident that the removal of content snapshot does in fact contribute somewhat to the Wiki join tasks. Numerical sketches contribute to ECB union which is perhaps because approximately 50% of the columns are not strings. Removal of MinHashes mirrors the result from earlier, which is, join tasks are most affected. Surprisingly, the removal of MinHashes affected the subset task; even though numerical sketches by themselves produced good performance. It is possible that there are complex interactions between the sketches. In summary, we answer Q2 from our ablation studies that different sketches play different roles across tasks, and may even interact to determining data discovery performance.
Table 3. Importance of different sketches on TabSketchFM performance for the different tasks. Performance of sketches that are equivalent (within 2 points) to overall performance is highlighted in green. TUS-SANTOS is not considered because it can be performed based on column headers alone.
Using only
Tasks MinHash Numerical Content TabSketchFM
sketches sketches snapshot (w/ everything)
Wiki Union (F1) 0.914 0.804 0.897 0.940
ECB Union (R2) 0.829 0.498 0.752 0.897
Wiki Jacc. (R2) 0.537 0.318 0.314 0.577
Wiki Cont. (R2) 0.628 0.252 0.301 0.587
Spider-Op. (F1) 0.831 0.817 0.797 0.831
ECB Join (F1) 0.874 0.812 0.815 0.856
CKAN Subset (F1) 0.431 0.984 0.431 0.986
Table 4. Removing a sketch to study the impact of it on TabSketchFM performance for the different tasks. Performance of sketches that declined more than 2 points compared to overall performance is highlighted in red.
Removing only
Tasks MinHash Numerical Content TabSketchFM
sketches sketches snapshot (w/ everything)
Wiki Union (F1) 0.933 0.927 0.931 0.940
ECB Union (R2) 0.770 0.872 0.897 0.897
Wiki Jacc. (R2) 0.425 0.565 0.519 0.577
Wiki Cont. (R2) 0.358 0.598 0.559 0.586
Spider-Op. (F1) 0.814 0.851 0.847 0.831
ECB Join (F1) 0.812 0.855 0.846 0.855
CKAN Subset (F1) 0.431 0.431 0.980 0.986
Further, in the technical report  [^10] Made available with other artifacts at https://anonymous.4open.science/r/tabsketchfm-6CDC/README.md we present a few more analysis to understand the robustness and sensitivity of TabSketchFM. Due to the lack of space, we only summarize a few important results here. First, we compare different hashing functions including: MMH3, FarmHash, SHA, Blake2S, xxhash and find that TabSketchFM is robust to the hashing implementations. Second, Pretraining is crucial to learn the data discovery tasks on smaller datasets. Third, large transformer-based models are crucial with sketches. Upon replacing the BERT model with MLP the performance is significantly impacted. Fourth, Minhash sketches support extremely large tables of enterprise data lakes. Using text tokens, shinglings, or Hashing Vectorizers is not scalable.
Figure 8. F1 on wiki join search for varying k.
Figure 9. F1 on Santos search for varying k.
Figure 10. F1 on TUS search for varying k.
6.3. Application to search
In the different search benchmarks we describe below, we combined the embeddings of columns derived from TabSketchFM models along with value embeddings for each column because as we describe in the introduction, using numerical and MinHash vectors as inputs to the transformer model made it difficult to include any cell values of a given column. To add cell values, which can contain important semantic information about a column, we created a single sentence by concatenating the top 100 values of a columnn as a sentence, and then embedded it using an off the shelf sentence embedding model (sentence-transformers/all-MiniLM-L12-v2). The embedding from TabSketchFM was concatenated to the value embeddings from the sentence embedder after normalizing them so the means and variances of the two vectors were in the same scale.
6.3.1. Join Search.
Here, the task is to find all the tables from the data lake that can be joined on one or more columns with the given query table. Note that prior join works (Zhu et al., 2019; Esmailoghli et al., 2022a; Dong et al., 2021) evaluate the efficiency of searching for data lake columns having value overlap with the query column(s), and assume that the columns with overlapping values are always joinable. However, we are interested in whether it makes sense to join tables. For example, two columns: (i) explaining people’s Ages with values such as (50, 24, 66, 78, 92, 95) and (ii) explaining students’ marks out of 100 (rounded to integer) with values such as (78, 92, 95, 100, 20) can have overlapping values but it is not sensible to join them because they have different semantics. Because TabSketchFM is built to generate contextual embeddings of columns that include value overlap, column headings, its context with other columns, and so on, it should be possible to get rid of such useless joins. To our knowledge, there are no such benchmarks with annotated gold standards that consider sensibility. So, we construct a search benchmark named Wiki Join from Wikidata, by following the cell-entity mapping approach used in the curation of the Wiki Jaccard dataset (see Section 5.2). Precisely, we generate the tables and then create a list of all the pairs of joinable columns, i.e., columns with the same entity annotations and overlapping values. We consider the columns having the same entity annotations to be sensible to join. We assign each pair an associated overlap score, which is the Jaccard score of the entity annotation sets. For each column in the list of pairs, we create a ranked list of joinable columns in descending order of their scores.
We used existing join search methods such as LSHForest and JOSIE (Zhu et al., 2019), along with embeddings derived from fine tuned versions of other pretrained models such as TABERT as our baselines. A recent embedding-based join search system, WarpGate (Cong et al., 2023), is not available publicly. So, we implemented a similar Glove-embedding-based baseline (Pennington et al., 2014) (EmbedJoin) that embeds columns into an embedding space and performs a search over HNSW index (Malkov and Yashunin, 2020). Other openly available systems, such as MATE (Esmailoghli et al., 2022b) were not easy to install due to software versioning requirements that we could not resolve. Figure 8 shows the significant advantage of TabSketchFM over all other baselines.
6.3.2. Union Search.
As in previous works (Fan et al., 2023; Hu et al., 2023), we use two publicly available benchmark datasets: $\mathbf{TUS_{Small}}$ and $\mathbf{SANTOS_{Small}}$ from  (Nargesian et al., 2018; Khatiwada et al., 2023a) because the larger versions of the dataset are not labeled. As a baseline, we compare against state-of-the-art table union search techniques: $\mathbf{D^{3}L}$ (Bogatu et al., 2020), SANTOS (Khatiwada et al., 2023a), and Starmie (Fan et al., 2023). We report results on $\mathbf{D^{3}L}$, SANTOS, and Starmie using their publicly available code. We could not compare against $AutoTUS$ (Hu et al., 2023), another recent union search system, because its code was not public at the time of writing this paper. For each technique, we use default parameters suggested in the respective papers. The results for TabSketchFM compared to the baselines show that TabSketchFM matched the best system (Fan et al., 2023) on SANTOS, see Figure 9. For TUS, on smaller K, STARMIE was better as shown in Figure 10, due to better recall.
6.3.3. Generalization across tasks and domains.
Figures 11, 12, and 13, show that the fine tuned cross encoder models generalize rather well across tasks as well as domain, which is encouraging particularly because it shows that the models can be trained on a datalake that is quite different from the datalake that it is deployed to. In enterprise contexts, this is an important requirement, because it means that we can basically train models offline and apply it online. All models shown in the Figure are models that include the value embeddings from a column for maximal generalization, but using the models without values also produced the same type of generalization.
Figure 11. F1 on wiki join transfer on search for varying k.
Figure 12. F1 on Santos transfer on search for varying k.
Figure 13. F1 on TUS transfer on search for varying k.
7. Conclusion and Discussion
We presented TabSketchFM, a novel transformer-based tabular representation model that inputs the table in the form of several sketches. We empirically show that cross encoders based on TabSketchFM outperform tabular models that serialize tables as texts, for three different data discovery tasks. We also show that the finetuned TabSketchFM can be effectively used to rerank the union and join search results. To further the research, we share 8 new datasets (LakeBench) that can be used to finetune neural models to understand tables for data discovery tasks. In the future, we will identify other data discovery tasks that TabSketchFM can be useful, and also make the model more robust for example, by incorporating other hashing mechanisms and cell values to create the sketches.